# Exploring a Patch-Wise Approach for Privacy-Preserving Fake ID Detection

### v1

@@ Abstract
In an increasingly digitalized world, verifying the authenticity of ID documents has become a critical challenge for real-life applications such as digital banking, crypto-exchanges, renting, etc. This study focuses on the topic of fake ID detection, covering several limitations in the field. In particular, no publicly available data from real ID documents exists, and most studies rely on proprietary in-house databases that are not available due to privacy reasons. In order to shed some light on this critical challenge that makes difficult to advance in the field, we explore a trade-off between privacy (i.e., amount of sensitive data available) and performance, proposing a novel patch-wise approach for privacy-preserving fake ID detection. Our proposed approach explores how privacy can be enhanced through: i) two levels of anonymization for an ID document (i.e., fully- and pseudo-anonymized), and ii) different patch size configurations, varying the amount of sensitive data visible in the patch image. Also, state-of-the-art methods such as Vision Transformers and Foundation Models are considered in the analysis. The experimental framework shows that, on an unseen database (DLC-2021), our proposal achieves 13.91% and 0% EERs at patch and ID document level, showing a good generalization to other databases. In addition to this exploration, another key contribution of our study is the release of the first publicly available database that contains 48,400 patches from both real and fake ID documents, along with the experimental framework and models, which will be available in our GitHub [^1] https://github.com/BiDAlab/ExploringFakeID-Patches.
@@ 1 Introduction
The veracity of digital content is one of the great challenges of society nowadays [2]. With the rapid advances made in the field of GenerativeAI [6], it is easy to synthesize non-existent content [15, 22], or to modify existing one [18], using simple toolbox available on the internet. Although these methods can be used for good purposes, e.g., correct biases or improve performance in some scenarios with limited data [7, 21], they can also be used for malicious purposes such as DeepFakes [24, 20] or misinformation [14]. In particular, the synthesis of non-existent fake IDs or the manipulation of real ones have started gaining attention world-wide due to their impressive realism [16, 23].
@@
Recent news have unveiled how fake ID documents are used nowadays for several frauds, such as under-age alcohol purchase [^2] https://www.nytimes.com/2025/02/13/nyregion/students-high-tech-fake-ids.html or impersonation for opening accounts in digital services such as crypto-exchanges or digital banks [^3] https://www.404media.co/inside-the-underground-site-where-ai-neural-networks-churns-out-fake-ids-onlyfake/, which use the Know Your Customer (KYC) verification system. This problem has been exacerbated by the rapid advance of Artificial Intelligence (AI) and GenerativeAI, being possible even to buy fake documents (ID, passport, driving license, etc.) on websites for a reasonable price (20\$), becoming a critical challenge in terms of security. Few studies have preliminary analyzed the problem of fake ID detection, proposing very valuable ideas and resources [3, 16, 10]. However, there are several limitations in the field that must still be covered to properly advance in this research line.
@@
On one hand, there are no public databases that contain any real (a.k.a. bona fide) data. This is mainly produced due to privacy concerns, as the information included in ID documents is very sensitive. Previous studies in the literature have always considered private in-house databases that are not publicly available [3, 16, 10]. This results in two main limitations: i) the lack of standard benchmarks to properly compare novel approaches with the state of the art in fake ID detection, and ii) the limited advance in the topic due to the lack of real data, as can be seen in recent international challenges carried out in IJCB 2024 conference [23].
@@
Figure 1: Graphical representation of our proposed patch-wise approach for privacy-preserving fake ID detection, exploring a trade-off between performance and privacy (i.e., amount of sensitive data available as input to the model).
@@
One of the first databases introduced in the literature that includes different types of fake IDs (a.k.a. presentation attacks) is the MIDV database family [1, 5, 4]. With the original purpose of Optical Character Recognition (OCR), the authors synthetically created a set of physical ID documents and passports. They used several digital templates from multiple countries which they filled with names and addresses from Wikipedia and artificially generated faces. The authors increased the number of fake documents with time as they released different versions of the databases. Later on, Polevoy et al. used in [19] the documents of the MIDV-family to create fake documents considering three different Presentation Attack Instrument (PAI) species: color print, gray print and screen. They also provided real samples, although they should be considered fake documents as they are not real, just prints in higher quality (i.e., glossy print). Recently, Benalcazar et al. proposed in [3] the use of Generative Adversarial Networks (GANs) to create synthetic Chilean ID cards. The proposed GAN was trained using only real Chilean ID samples. Although the data synthesized by the authors is not strictly real ID samples, this was presented as a good idea to partially cover the lack of real data, e.g., as a data augmentation strategy. In [10], Gonzalez and Tapia proposed a more sophisticated fake ID generation, printing over a poly-carbonate card, which resembles even more the appearance of real ID documents.
@@
All the variability mentioned in terms of digital/physical attacks results in very poor fake ID detection performances, especially in unconstrained scenarios. One of the first fake ID detection approaches was presented in [16], where the authors trained a Convolutional Neural Network (CNN) to perform pixel-wise classification to detect different types of PAI s. The authors created their own private in-house database with both real and fake documents including print and screen PAIs. Gonzalez and Tapia proposed in [10] a two-stage system which first uses a neural network to evaluate if the submitted ID card is real or fake, concretely using digital PAIs of type composite or synthetic. The second network classified the submitted ID document between real or fake, concretely physical PAIs such as print, screen or PVC.
@@
Finally, we would like to highlight the recent fake ID competition celebrated in IJCB 2024 [23]. Due to privacy concerns, the organizers did not provide participants with a database for training the proposed fake ID detectors, stressing the challenge of the task. The evaluation of the submitted detectors was done by the organizers using a sequestered private database. Regarding the performance, very poor results were achieved by all teams in the competition, where the winner of the competition achieved an Equal Error Rate (EER) of 21.87%. This result highlights the importance to promote this challenging research line.
(a) Non-anonymized ID document.
(b) Pseudo-Anonymized ID document.
(c) Fully-Anonymized ID document.
@@
Figure 2: Graphical examples of the three different anonymization configurations considered in our proposed method, with different levels of privacy in terms of the sensitivity of the information available in the ID document.
@@
In this paper, we propose to tackle the task of fake ID detection from a different perspective, exploring a novel trade-off between performance and privacy. Concretely, instead of considering the whole ID document as input to the models, our proposed approach explores how privacy can be enhanced through: i) two levels of anonymization for an ID document (i.e., fully- and pseudo-anonymized), and ii) different patch size configurations, varying the amount of sensitive data visible in the patch image. These patches are then introduced to a deep learning model, classifying them as real or fake. Fig. 1 provides a representation of our proposed approach. Our main contributions are as follows:
@@
  • We propose a novel patch-wise approach for privacy-preserving fake ID detection, exploring several configurations in terms of performance and privacy. In particular, as can be seen in Fig. 1, we consider three different scenarios in terms of anonymization (non-, pseudo-, and full-anonymized ID) and also in terms of the size of the patches ($128\times 128$, $64\times 64$, and $32\times 32$), providing different levels of privacy and amount of training samples depending on the particular scenario.
@@
  • We provide the first publicly available database that contains patches from both real and fake ID documents. Concretely, the proposed database contains 30 real ID documents from 30 different subjects. In addition, for each real ID, different methods have been used in order to produce high-quality fake IDs, considering different PAI species (i.e., print and screen). The pseudo- and fully-anonymized ID data with the best patch size configuration ($64\times 64$) will be available in our GitHub [^4] https://github.com/BiDAlab/ExploringFakeID-Patches for research purposes.
@@
  • In addition, we explore different state-of-the-art deep learning methods such as Residual Network (ResNet) [11], Vision Transformer (ViT) [9], or DINOv2 Foundation Model [17] for the extraction of discriminative real/fake patches, considering pre-trained models and fine-tuning. The best fake detectors will be available in our GitHub repository for research purposes.
@@
The rest of the paper is organized as follows. In Sec. 2 we explain the details of our proposed patch-wise approach for privacy-preserving fake ID detection. Sec. 3 provides the acquisition details of our novel fake ID database. In Sec. 4, we explain the experimental framework of the study, including the experimental protocol and evaluation metrics. Sec. 5 shows the results achieved by our proposed method. Finally, the key conclusions are included in Sec. 6.
2 Proposed Method
@@
Fig. 1 provides a graphical representation of our proposed patch-wise approach for privacy-preserving fake ID detection. As can be seen, it comprises two main modules: i) a privacy-preserving patch extractor, and ii) a fake patch detector. Finally, a score fusion of the individual patches is considered to detect a whole ID document as real or fake. We provide next all the details.
@@ 2.1 Privacy-Preserving Patch Extractor
Given an ID document, we first explore three different scenarios in terms of anonymization, considering different configurations in terms of privacy, as can be seen in Fig. 2:
@@
    Non-Anonymized ID: there is no anonymization. All information included in the ID document is available to detect real/fake IDs.
@@
    Pseudo-Anonymized ID: sensitive fields are partially anonymized, such as the expiration date, ID document number, the face of the owner, and the support number, leaving some sections of information available. The name and surname are never displayed together (i.e., one of them is always anonymized). In addition, they are partially occluded showing only some random characters so that it is impossible to know the full name/surname of the owner.
@@
    Fully-Anonymized ID: all fields with sensitive information of the owner (i.e., text, face image, and handwritten signature) are completely anonymized and only patches from outside of the sensitive information are gathered.
In addition to the anonymization scenario, we explore the extraction of patches with different sizes: $128\times 128$, $64\times 64$, and $32\times 32$ pixels. The main motivation for this is to also increase privacy, showing less sensitive information of the subject as the image size of the patch is reduced. Graphical examples of real and fake patches at different sizes can be seen in Fig. 3, where we can see that the patches alone do not contain any sensitive information.
@@ 2.2 Fake Patch Detector
After the extraction of patches from an ID document, we train three different state-of-the-art deep learning models to classify each patch between real or fake.
@@
  • ViT-B/16 Vision Transformer [9]: the ViT was the first architecture that used the self-attention layers introduced in the Transformer [25] for image classification. Concretely, we consider the ViT-based model pre-trained on the ImageNet database [8].
  • DINOv2 [17]: this model uses the ViT architecture to learn visual feature representations, considering a self-supervised approach based on distillation [12]. This is based on a teacher-student scheme with a slight modification. While the student is fed with both the whole image and local patches of said image, the teacher model only received the whole image. The core idea of this training procedure is for the teacher model to try to match the embeddings distribution of the student. This patch strategy, combined with centering and momentum updates for the teacher model weights, helped DINO learn semantic representations without labels.
@@
The selected deep learning models have been modified in order to adapt them to our problem, fake ID detection. First, for both ResNet-18 and ViT-B/16 models pre-trained with ImageNet database, we replaced the final softmax layer (1,000 classes) with a fully-connected layer based on a sigmoid activation in order to provide a continuous score for real (values close to 0) and fake (values close to 1) IDs. Regarding DINOv2, as it was trained using an unsupervised approach with the purpose of extracting visual features, we only added a fully-connected layer with a sigmoid activation. Regarding the fine-tuning , we decided to keep the backbones of the models frozen and only train the new fully-connected layers, given the limited amount of training data. For any patch size, the input was resized to $224\times 224$, which is the input shape of all three models. The loss function selected for training these models was Binary Cross Entropy (BCE). The selected optimizer was Adam [13] with a learning rate of $\alpha=0.00015$ and exponential decays for the momentum estimators of $\beta_{1}=0.9$ and $\beta_{2}=0.999$. All the models were trained for 150 epochs, but with an early-stopping condition that ended the training if no improvements were obtained in the loss value for the validation split after 12 epochs.
Finally, although our proposed method is designed to detect each patch as real or fake, it also allows for whole ID document classification. This is referred throughout the paper as “Patch Level” or “ID Document Level”. Concretely, for the “ID Document Level”, in the present study we consider a score fusion of the individual patches based on the mean of the predictions, obtaining a final score between 0 and 1. Other approaches such as majority voting could be used for the fusion, although this type of mechanism would only provide a binary output (not continuous value) with less discriminative information of the decision.
3 Proposed Database
@@ 3.1 Real and Fake IDs: Acquisition
Due to the lack of publicly available databases, one of the key contributions of our study is the acquisition and release of a new database, comprising both real and fake ID documents. For the real IDs, there are a total of 30 images of Spanish ID documents, each one belonging to a different subject (i.e., 30 subjects in total). Spanish electronic ID documents appearance have changed over the years, which was also taken into account while capturing the data, considering three different versions/templates in the database. Referring to the capturing device, we used a Redmi 9C NFC smartphone, a low-end device with a sensor of 13 MP, a wide $f$/2.2 aperture, that captures images in 4K resolution and 4:3 aspect with High Dynamic Range (HDR). All pictures were saved in a JPEG format. Pictures were taken with a vertical distance of approximately 15 centimeters from the ID document.
Regarding the fake ID documents, we consider two types of PAIs: print and screen. In both cases, in order to generate high-quality fake IDs, we first scanned the real ID documents using a HP ScanJet 8270 scanner at 600 Dots Per Inch (DPI). After getting the corresponding digital copies for each ID document, we created the print PAI following a similar approach as in [1]. We used an EPSON ET-2850 Wi-Fi and regular paper to print the scanned versions of the real IDs, which then were laminated to improve the realism using a Fellowes Lunar A4 thermal laminator. Regarding the screen PAI, we displayed the ID images on a MacBook Pro 14” XDR display, with a resolution of $3024\times 1964$, in full screen mode and took the pictures so that the whole displayed document covered the whole camera preview of the smartphone. We selected this screen as it was the panel with more Pixels Per Inch (PPI) available (i.e., 254 PPI). This aspect is very important for high-quality fake IDs as the space between pixels in a screen is one of the most evident features of screen PAI [19]. Additionally, we took special care in avoiding any aliasing or interference patterns, such as moiré patterns when taking the pictures. The proposed acquisition resulted in 90 real/fake IDs in total: 30 real IDs, 30 print fake IDs and 30 screen fake IDs.

### v2

# FakeIDet: Exploring Patches for Privacy-Preserving Fake ID Detection

@@ Abstract
Verifying the authenticity of identity documents (IDs) has become a critical challenge for real-life applications such as digital banking, crypto-exchanges, renting, etc. This study focuses on the topic of fake ID detection, covering several limitations in the field. In particular, there are no publicly available data from real IDs for proper research in this area, and most published studies rely on proprietary internal databases that are not available for privacy reasons. In order to advance this critical challenge of real data scarcity that makes it so difficult to advance the technology of machine learning-based fake ID detection, we introduce a new patch-based methodology that trades off privacy and performance, and propose a novel patch-wise approach for privacy-aware fake ID detection: FakeIDet. In our experiments, we explore: i) two levels of anonymization for an ID (i.e., fully- and pseudo-anonymized), and ii) different patch size configurations, varying the amount of sensitive data visible in the patch image. State-of-the-art methods, such as vision transformers and foundation models, are considered as backbones. Our results show that, on an unseen database (DLC-2021), our proposal for fake ID detection achieves 13.91% and 0% EERs at the patch and the whole ID level, showing a good generalization to other databases. In addition to the path-based methodology introduced and the new FakeIDet method based on it, another key contribution of our article is the release of the first publicly available database that contains 48,400 patches from real and fake IDs, called FakeIDet-db, together with the experimental framework [^1] https://github.com/BiDAlab/FakeIDet-db.
@@ 1 Introduction
Verification of the veracity of digital content is one of the great challenges facing society today [2]. With the rapid advances made in the field of Generative AI [6], it is easy to synthesize non-existent content [18, 28], or to modify existing one [23]. Although these methods can be used for good purposes, e.g., correct biases [19, 24] or improve performance in some scenarios with limited data [8, 11, 27], they can also be used for malicious purposes such as generating injurious DeepFakes [26, 30] or misinformation [17]. In particular, the synthesis of non-existent fake IDs or the manipulation of real ones has started gaining attention due to their impressive realism [20, 29]. This is strongly related to the area of presentation attacks in biometrics [7, 11, 14].
@@
Recent news has revealed how fake identity documents (IDs) are used nowadays for several frauds, such as underage purchases of alcohol [^2] https://www.nytimes.com/2025/02/13/nyregion/students-high-tech-fake-ids.html or impersonating to open accounts in digital services such as crypto-exchanges or digital banks [^3] https://www.404media.co/inside-the-underground-site-where-ai-neural-networks-churns-out-fake-ids-onlyfake/, which use Know Your Customer (KYC) verification systems. This problem has been exacerbated by the rapid advance of Artificial Intelligence (AI) and Generative AI, which have allowed one to buy fake ID documents (identity card, passport, driving license, etc.) on websites for a reasonable price (20\$), making this a critical challenge in terms of security. Few studies have preliminarily analyzed the problem of fake ID detection, proposing very valuable ideas and resources [3, 12, 20]. However, there are several limitations in the field that need to be covered to properly advance this research line.
@@
On the one hand, there are no public databases that contain real (a.k.a. bona fide) data. This is mainly produced due to privacy concerns, as the information included in IDs is very sensitive. Previous studies in the literature have always considered private in-house databases not publicly available [3, 12, 20]. This results in two main limitations: i) the lack of standard benchmarks to properly compare novel approaches with the state of the art in fake ID detection, and ii) the poor performance of existing fake ID detectors due to the lack of official real data for training machine learning-based approaches, as can be seen in a recent international challenge carried out at the IJCB 2024 conference [29].
@@
Figure 1: Graphical representation of FakeIDet, our proposed patch-wise approach for privacy-preserving fake ID detection, exploring a trade-off between performance and privacy (i.e., amount of sensitive data available as input to the model).
@@
One of the first databases introduced in the literature that includes different types of fake IDs is the MIDV database family [1, 5, 4]. With the original purpose of Optical Character Recognition (OCR), the authors synthetically created a set of physical IDs and passports. They used several digital templates from multiple countries that they filled with names and addresses from Wikipedia and artificially generated faces. The authors increased the number of fake IDs with time as they released different versions of the databases. Later, Polevoy et al. used in [25] the documents of the MIDV-family to create fake documents considering three different Presentation Attack Instrument (PAI) species: color print, gray print, and screen [14]. They also provided real samples, although they should be considered fake documents, as they are not real, just prints in higher quality (i.e., glossy print). Recently, Benalcazar et al. proposed in [3] the use of Generative Adversarial Networks (GANs) to create synthetic Chilean IDs. The proposed GAN was trained using only real Chilean ID samples. Although the data synthesized by the authors are not strictly real ID samples (note that GAN traces can be used to spot these synthetic samples [21]), this was presented as a good idea to partially cover the lack of real data, for example, as a data augmentation strategy. In [12], Gonzalez and Tapia proposed a more sophisticated fake ID generation, printing on a Poly-Vinyl Chloride (PVC) card, which resembles even more the appearance of real IDs.
@@
All the variability mentioned in terms of digital/physical attacks results in very poor fake ID detection performances, especially in unconstrained scenarios. One of the first fake ID detection approaches was presented in [20], where the authors trained a Convolutional Neural Network (CNN) to perform pixel-wise classification to detect different types of PAI s. The authors created their own private in-house database with both real and fake documents, including print and screen PAIs. Gonzalez and Tapia proposed in [12] a two-stage system which first uses a neural network to evaluate if the submitted ID card is real or fake, concretely using digital PAIs of type composite or synthetic. The second network classified the ID considered between real or fake for physical PAIs such as printing, screen or PVC.
@@
Finally, we would like to highlight the recent fake ID contest celebrated in IJCB 2024 [29]. Due to privacy concerns, organizers did not provide participants with a database to train the proposed fake ID detectors, stressing the difficulty of the challenge. The evaluation of the detectors submitted was done by the organizers using a private sequestered database. Regarding performance, very poor results were achieved by all teams in the competition, where the winner of the competition achieved an Equal Error Rate (EER) of 21.87%. This result highlights how challenging fake ID detection is in realistic settings.
(a) Non-anonymized ID.
(b) Pseudo-Anonymized ID.
(c) Fully-Anonymized ID.
@@
Figure 2: Graphical examples of the three different anonymization configurations considered in our proposed method, with different levels of privacy in terms of the sensitivity of the information available in the ID.
@@
In this paper, we propose to tackle the task of fake ID detection from a different perspective, exploring a novel trade-off between performance and privacy. Concretely, instead of considering the whole ID as input to the models, our proposed approach explores how privacy can be enhanced through: i) two levels of anonymization for an ID (i.e., fully- and pseudo-anonymized), and ii) different patch size configurations, varying the amount of sensitive data visible in the patch image. These patches are then introduced into a deep learning model, classifying them as real or fake. Fig. 1 provides a representation of our proposed approach. Our main contributions are as follows:
@@
  • We propose a novel patch-wise approach for privacy-preserving fake ID detection, called FakeIDet, exploring several configurations in terms of performance and privacy. In particular, as can be seen in Fig. 1, we consider three different scenarios in terms of anonymization (non-, pseudo-, and fully-anonymized ID) and also in terms of the size of the patches ($\times 128$ 128, $\times 64$ 64, and $\times 32$ 32), providing different levels of privacy and amount of training samples depending on the particular scenario.
@@
  • We provide the first publicly available database that contains patches from both real and fake IDs, called FakeIDet-db. The proposed database contains 30 real IDs from 30 different subjects. In addition, for each real ID, different methods have been used in order to produce high-quality fake IDs, considering different PAI species (i.e., print and screen). The pseudo- and fully-anonymized ID data with the best patch size configuration ($\times 64$ 64) will be available on our GitHub [^4] https://github.com/BiDAlab/FakeIDet-db for research purposes.
@@
  • In addition, we explore different state-of-the-art deep learning methods such as Residual Network (ResNet) [13], Vision Transformer (ViT) [10], or DINOv2 Foundation Model [22] for the extraction of discriminative real/fake patches, considering pre-trained models and fine-tuning.
@@
The remainder of the paper is organized as follows. In Sec. 2 we explain the details of our proposed patch-wise approach for privacy-preserving fake ID detection, FakeIDet. Sec. 3 provides the acquisition details of our new fake ID database, FakeIDet-db. In Sec. 4, we explain the experimental framework of the study, including the experimental protocol and evaluation metrics. Sec. 5 shows the results achieved by our proposed method. Finally, the conclusions are presented in Sec. 6.
2 FakeIDet: Proposed Method
@@
Fig. 1 provides a graphical representation of FakeIDet, our proposed patch-wise approach for privacy-preserving fake ID detection. As can be seen, it comprises two main modules: i) a privacy-preserving patch extractor, and ii) a fake patch detector. Finally, a score fusion of the individual patches is considered to detect a whole ID as real or fake. We provide next all the details.
@@ 2.1 Privacy-Preserving Patch Extractor
Given an ID, we first explore three different scenarios in terms of anonymization, considering different configurations in terms of privacy, as can be seen in Fig. 2:
@@
    Non-Anonymized ID: there is no anonymization. All information included in the ID is available to detect real/fake IDs.
@@
    Pseudo-Anonymized ID: sensitive fields are partially anonymized, such as the expiration date, ID number, the face of the owner and the support number, leaving some sections of information available. The name and surname are never displayed together (i.e., one of them is always anonymized). In addition, they are partially occluded, showing only some random characters, so that it is impossible to know the full name/surname of the owner.
@@
    Fully-Anonymized ID: all fields with sensitive information about the owner (i.e., text, face image, and handwritten signature) are completely anonymized and only patches from outside of the sensitive information are collected.
In addition to the anonymization scenario, we explore the extraction of patches of different sizes: $\times 128$ 128, $\times 64$ 64, and $\times 32$ 32 pixels. The main motivation for this is also to increase privacy, showing less sensitive information of the subject as the image size of the patch is reduced. Graphical examples of real and fake patches of different sizes can be seen in Fig. 3.
@@ 2.2 Fake Patch Detector
After the extraction of patches from an ID, we train three different state-of-the-art deep learning models to classify each patch between real and fake.
@@
  • ViT-B/16 Vision Transformer [10]: the ViT was the first architecture to use the self-attention layers introduced in Transformers [31] for image classification. Concretely, we consider the ViT-based model pre-trained in the ImageNet database [9].
  • DINOv2 [22]: this model uses the ViT architecture to learn visual feature representations, considering a self-supervised approach based on distillation [15]. This is based on a teacher-student scheme with a slight modification. While the student is fed with the whole image and local patches of it, the teacher model received only the whole image. The core idea of this training procedure is that the teacher model tries to match the embedding distribution of the student. This patch strategy, combined with centering and momentum updates for teacher model weights, helped DINO learn semantic representations without labels.
@@
The selected deep learning models have been modified in order to adapt them to our problem, fake ID detection. First, for both ResNet-18 and ViT-B/16 models pre-trained with ImageNet database, we replaced the final softmax layer (1,000 classes) with a fully connected layer based on a sigmoid activation in order to provide a continuous score for real (values close to 0) and fake (values close to 1) IDs. Regarding DINOv2, as it was trained using an unsupervised approach with the purpose of extracting visual features, we only added a fully connected layer with a sigmoid activation. Regarding the fine-tuning, we decided to keep the backbones of the models frozen and only train the new fully connected layers, given the limited amount of training data. For any patch size, the input was resized to $\times 224$ 224, which is the input shape of the three models. The loss function selected to train these models was Binary Cross Entropy (BCE). The selected optimizer was Adam [16] with a learning rate of $\alpha=0.00015$ and exponential decays for the momentum estimators of $\beta_{1}=0.9$ and $\beta_{2}=0.999$. All models were trained for 150 epochs, but with an early stopping condition that ended the training if no improvements were obtained in the loss value for the validation split after 12 epochs.
Finally, although our proposed method is designed to detect each patch as real or fake, it also allows for whole ID classification. This is referred to throughout the paper as “Patch Level” or “ID Level”. Concretely, for the “ID Level”, in the present study, we consider a score fusion of the individual patches based on the mean of the predictions, obtaining a final score between 0 and 1. Other approaches such as majority voting could be used for the fusion, although this type of mechanism would only provide a binary output (not continuous value) with less discriminative information of the decision.
3 FakeIDet-db: Proposed Database
@@ 3.1 Real and Fake IDs: Acquisition
Due to the lack of publicly available databases, one of the key contributions of our study is the acquisition and publication of a new database, FakeIDet-db, comprising real and fake identification documents. For real IDs, there are a total of 30 images of Spanish IDs, each one belonging to a different subject (that is, 30 subjects in total). Spanish electronic IDs have changed over the years, which was also taken into account when capturing the data, considering three different versions/templates in the database. Referring to the capture device, we used a Redmi 9C NFC smartphone, a low-end device with a sensor of 13 MP, a wide $f$/2.2 aperture, which captures images in 4K resolution and 4:3 aspect with High Dynamic Range (HDR). All pictures were saved in JPEG format (very high quality, so compression artifacts are not noticeable, see, e.g., Fig. 3). The photos were taken with a vertical distance of approximately 15 centimeters from the ID.
Regarding fake IDs, we consider two types of PAIs: print and screen. In both cases, in order to generate high-quality fake IDs, we first scanned the real IDs using an HP ScanJet 8270 scanner at 600 Dots Per Inch (DPI). After getting the corresponding digital copies for each ID, we create the print PAI following a similar approach to [1]. We used an EPSON ET-2850 Wi-Fi and regular paper to print the scanned versions of the real IDs, which then were laminated to improve the realism using a Fellowes Lunar A4 thermal laminator. Regarding the screen PAI, we display the ID images on a MacBook Pro 14” XDR display, with a resolution of 3024×19643024$\times$ 19643024 × 1964, in full-screen mode and take the pictures so that the whole document displayed covered the whole camera preview of the smartphone. We selected this screen because it was the panel with more Pixels Per Inch (PPI) available (that is, 254 PPI). This aspect is very important for high-quality fake IDs as the space between pixels in a screen is one of the most evident features of the screen PAI [25]. In addition, we took special care to avoid any aliasing or interference patterns, such as Moiré patterns, when taking pictures. The proposed acquisition resulted in 90 real/fake IDs in total: 30 real IDs, 30 print fake IDs, and 30 screen fake IDs.